Henry pizzas for a party and organized the information into a table.
****************************************************************************

Type of Pizza Number of Pizzas Price of One Pizza
-----------------------------------------------------------------------------------------------------
Large Cheese. X. $13.95
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Medium Pepperoni. 2. $ 11.95
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

If Henry paid a total of 93.65, how many large pizzas did he order?



Answer by Mini-(17)

X = ( 93.65 - ( 11.95 * 2 ) ) / 13.95
X = ( 93.65 - 23.90 ) / 13.95
X = 69.75 / 13.95
X = 5